OpenStudy (anonymous):

1 Create a formula that will tell you how many people are in the US if 1/4th of the people equals 57.7 million

OpenStudy (anonymous):

no. of people = 4 x 57,700,000

OpenStudy (anonymous):

\[\large \frac{1}{4}=\frac{57700000}{x}\] \[\large x=4(57,700,000)\]

OpenStudy (anonymous):

so the answer is

OpenStudy (anonymous):

Multiply for your answer....